OpenStudy (anonymous):

1. using the definition show that lim n/2n-3 =1/2 (the limit is as n approches to infinity).

OpenStudy (anonymous):

definition means that given any \(\epsilon >0\) there is an \(N_\epsilon\) such that if \(n>N\) you have \[|\frac{n}{2n-3}-\frac{1}{2}|<\epsilon\]

OpenStudy (anonymous):

you work backwards to find the \(N\) first algebra inside the absolute value signs, then it will be clear what \(N\) you need to pick, depending on \(\epsilon\)

OpenStudy (anonymous):

u r right, it is using epsilon..

OpenStudy (anonymous):

lol abusive because i did not show the necessary steps? diy

OpenStudy (anonymous):

1. using the definition show that lim n/2n-3 =1/2 (the limit is as n approches to infinity).

OpenStudy (anonymous):

\[\lim_{n \rightarrow \infty} \frac{ n }{ 2n-3 }\] Dividing Nr. and Dr. by n we find \[= \lim_{n \rightarrow \infty} \frac{ 1 }{ 2-3(\frac{ 1 }{ n })}\] \[n \rightarrow \infty, \frac{ 1 }{ n }\rightarrow 0\] Therefore \[= \frac{ 1 }{ 2-0}\] \[= \frac{ 1 }{ 2}\] Thus proved

OpenStudy (anonymous):

definition means that given any ϵ>0 there is an N ϵ such that if n>N you have

OpenStudy (anonymous):

using the definition show that lim n/2n-3 =1/2 (the limit is as n approaches to infinity).definition means that given any ϵ>0 there is an N ϵ such that if n>N

OpenStudy (anonymous):

show that of summation of 1/n2^n from 1 to infinity = ln2